OpenStudy (anonymous):

In a study of 235 adults, the mean heart rate was 82 beats per minute. Assume the population of heart rates is known to be approximately normal with a standard deviation of 5 beats per minute. What is the 98% confidence interval for the mean beats per minute? 80.9 – 86.3 70.9 – 83.3 81.2 – 82.8 80.9 – 83.3

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

The \((1-\alpha)\times100\%\) confidence interval has the form \[\left(\bar{x}-Z_{\alpha/2}\sqrt{\frac{\sigma^2}{n}},~\bar{x}+Z_{\alpha/2}\sqrt{\frac{\sigma^2}{n}}\right)\] where \(\bar{x}\) is the estimated population mean (\(\bar{x}=82\)), \(Z_{\alpha/2}\) is the critical \(z\) value for 98% confidence \(\left(Z_{\alpha/2}=Z_{.02/2}=Z_{.01}\approx2.33\right)\), \(\sigma\) is the sample standard deviation (\(\sigma=5\)), and \(n\) is the sample size (\(n=235\)). \[\left(82-2.33\frac{5}{\sqrt{235}},~82+2.33\frac{5}{\sqrt{235}}\right)=\left(81.24,~82.76\right)\]